OpenStudy (anonymous):

The perimeter of a regular hexagon is 36 and each side is represented by 2x . What is x?

OpenStudy (anonymous):

36+2=38 so you subtract 180 from 38 and you get your answer

OpenStudy (anonymous):

hexagon means 6 sides. Regular means the sides are all equal in length. perimeter = 2x+2x+2x+2x+2x+2x = 6 times 2x = 36 (which was given) (6)(2x) = 12 x = 36 therefore x = 36/12 = 3.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np